OpenStudy (anonymous):

The height of a triangle is 6 m more than its base. The area of the triangle is 56 m^2. Find the length of the triangle. A. 7m B. 8m C. 14m D. 15m

Mehek (mehek14):

do you know the formula for finding the area of a triangle?

OpenStudy (anonymous):

I honestly dont know this is a question from my sister shes in highschool

Mehek (mehek14):

A=(1/2)BH

Mehek (mehek14):

so you have 1/2(6+h)*h= 56

OpenStudy (anonymous):

Ok next

Mehek (mehek14):

bh = 112

Mehek (mehek14):

\(\Large{A = (1/2)bh = 56m^2}\) \(\Large{bh = 112}\) \(\Large{b(b+6m) = 112m^2}\)

OpenStudy (anonymous):

Ok which makes the answer 14 m correct?

Mehek (mehek14):

hold on I'm getting confused

Mehek (mehek14):

yes i believe it is 14

OpenStudy (anonymous):

Ok got it thx

Mehek (mehek14):

yw